The Advanced Photon Source
a U.S. Department of Energy Office of Science User Facility

APS-EFOG Meeting Minutes - 07/10/2017

Attendees: Bruce Glagola, Bruno Fieramosca, Nena Moonier, Clay White, Shane Flood, Dean Wyncott, and Patti Pedergnana

Bruce Glagola

  • The User Office requested that we check the user’s training when posting ESAFs.  Bruce reminded us to continue doing this.
  • Steve Davey is going on vacation from 7/18-7/30.  He asked for Patti to be back-up BSDRSC chair while he is out.  Bruce suggested that Julie Cross be a reference if this happens.  Davey will discuss this with John Connolly.
  • The group discussed the Floor Coordinator applicants.  Bruce will arrange a meeting with our HR representative and start scheduling interviews with the selected applicants.

Bruno Fieramosca

  • Bruno distributed photos of the 4-ID-C end-station showing an end cap.  We will now RSS tag the X-05 end-table and the end cap if these are installed in the future.  Patti will review the 4-ID and 29-ID procedures.
  • The liquid helium recovery project in 431/432 truck lock may take up to 3 weeks to complete.  John Freeland asked if they could wait to do some of the work during today’s machine studies as construction equipment was in the way of beamline access.
  • Bruno asked that when we issue an O2 monitor to also hand out the door placard.

Nena Moonier

  • Sectors 13, 14 and 15 had their ES&H walk-throughs today.  Sectors 16, 17, and 19 will be walked-through on Monday, 7/17/2017.

Patti Pedergnana

  • 16-BM-D’s door was adjusted today and a HP type C survey is needed tomorrow.
  • 16-ID-D’s radioactive sample experiment was broken down this morning.
  • Clay had Ben Reinhart move a methane cylinder from Sector 13 to the 12-BM cylinder cage in the gas yard.

Shane Flood

  • The water leak at the 23-ID Front End is in the corrector.  The corrector was taken out of service for now.  The repair will not be done until next week.
  • 27-ID had trouble opening their hutch door over the weekend.  Tim Clute and Jack Burke replaced a pneumatics box solenoid on the B door and it was working again.